# Manifolds from Partitions

### 15 Jan 2024 v1

@@ Abstract.
If $f$ maps a discrete $d$-manifold $G$ onto a $(k+1)$-partite complex $P$ then $H(G,f,P)$, the set of simplices $x$ in $G$ such that $f(x)$ contains at least one facet in $P$ is a $(d-k)$-manifold.
@@ 1. The theorem
$H(G,f,P)$ is a $(d-k)$-manifold if $f$ maps a $d$-manifold $G$ P-onto a $k$-complex $P\in\mathcal{P}$,
@@
Theorem (1) can be seen as a discrete inverse function theorem. It is the discrete analog of the fact that if $f:M\to N$ is a smooth map from a d-manifold $M$ to a $k$-manifold $N$ and $p$ is a regular value in $N$, then $f^{-1}(p)$ is a $d-k$-manifold. The discrete Sard result [7] is a special case, where we used a particular partition complex (topology) on the finite target set $\{-1,1\}^{k}$ of ${\rm sign}(f-c)$ given by a map $f:G\to\mathbb{R}^{k}$. Theorem (1) implies in particular that if $f:G\to y=\{0,\dots,k\}$ is surjective, then the set of simplices $x$ with $f(x)=y$ is a $(d-k)$-manifold. In [5], we looked maps $f:G\to\mathbb{R}$ which leads for $c$ not in $f(G)$ to $x\to{\rm sign}(x)\in\{-1,1\}$. If $f:G\to\mathbb{R}$ has no root and takes both positive and negative values, then $\{f=0\}$ is always a hyper-surface. This was already magic in the very special case $P=K_{2}$.
Figure 1. For Sard $G\to\mathbb{R}$, where studying $\{f=c\}$ is determined by the map ${\rm sign}(f-c)$ onto $\{(-1,1)\}$, the complex $K_{2}=K_{(1,1)}$ is the only relevant choice [5]. For Sard $G\to\mathbb{R}^{2}$ where ${\rm sign}(f-c)$ map into $\{(-1,1),(1,1),(-1,-1),(-1,-1)\}$, we needed to pick the unique 2-dim partition complex $P=K_{(1,1,2)}$ that exists for $n=4$ [7]. For $G\to\mathbb{R}^{3}$, the map ${\rm sign}(f-c)$ targets $\{-1,1\}^{3}$ which has 8 points. The 3-dimensional partitions of $n=8$ are $(5,1,1,1),(4,2,1,1),(3,3,1,1),(3,2,2,1),(2,2,2,2)$. In [7], we picked $p=(1,1,1,5)$ by forced the target signs to be $(-1,1,1),(1,-1,1),(1,1,-1)$ and requiring to reach a 4th point. This complex $P$ is displayed to the right. For maps into $\mathbb{R}^{m}$ we chose $p=(1,1,\dots,1,2^{m}-m))$ in [7].
@@
An interesting case for $G\to\mathbb{R}^{3}$ is when the target complex is $P=K_{(2,2,2,2)}$ which is a 3-sphere, the 3-dimensional cross polytop, one of the regular polytops in ambient dimension $4$. We can implement the quaternion group $Q$ on $P$, as we can identify the $8$ points in $P$ with $1,-1,i,-i,j,-j,k,-k$. In other words, a map $f$ from $G$ to the quaternion group $Q$ which not have an Abelian image, produces a co-dimension $3$ manifold $H(G,f,P)$, where $P$ is the 3-sphere simplicial complex structure on $Q$ coming from the partition $p=(2,2,2,2)$. In particular, any map from a 4-manifold to $Q$ with this property produces a co-dimension $3$ manifold $H(G,f,P)$ which must be a finite collection of 1-dimensional cycle graphs.
@@
If $G$ is a $d$-manifold and $f:G\to\{0,1,\dots,k\}$ is surjective then $H$ is a $(d-k)$-manifold.
@@
If $P$ is a k-dimensional partition complex and $y$ is a facet in $P$ which is in the image of $f$, then $f^{-1}(y)$ is a $(d-k)$-manifold with boundary. This allows us to construct without much effort manifolds with boundary or pictures of cobordism examples: if the boundary of the manifold is split into two sets $A,B$, then $A,B$ are by definition cobordant. In illustrations done before, like for the expository article [4] in 2012, we needed to construct by hand the ”hose manifold”. We later would use triangulations which numerical methods would provide when parametrizing surfaces $\vec{r}(u,v)$ where $(u,v)\subset\mathbb{R}^{2}$ is a closed region in the parameter plane. The code provided in the code section is shorter than what one needs to get in order to rely on numerical schemes used by computer algebra systems for drawing surfaces. The code works for manifolds of arbitrary dimension.
@@
The $(d-k)$-manifold $H(G,f,P)$ is the union of patches $f^{-1}(y)$, each of which is a $(d-k)$-manifold with boundary.
@@ 2. Partitions
Figure 4. All 15 integer partitions for $n=7$ and the corresponding 15 complexes. We used Cuisenaire tools in first grade [CuisenaireGattegno] were partitions were studied for the competition [3]. As far as I know, this was the first use of such a pedagogical tool for research rather than for teaching. Beside known results like the Euler pentagonal theorem, unusual results appeared like like $p(n)=\sum_{k=1}^{n}\sigma(k)p(n-k)/n$, where $\sigma(k)$ is the total number of prime factors of $k$. This can be seen by looking at $np(n)$ as the area of the Cuisenaire rectangle listing all the partitions.
@@
Figure 6. A 2-manifold obtained by taking the target complex $P=K_{(2,2,3)}$ and $G=I\oplus C_{13}$, where $I$ is the icosahedron complex. While the graph $G$ has only $12+13=25$ vertices, the graph $H(G,f,P)$ is much larger than $G$ as it is a subgraph of the Barycentric refinement of $G$.
@@
Lets look at a geometric characterization of the partition monoid $(\mathcal{P},\oplus)$. Let $\mathcal{Q}$ be the set of graphs which have the property that it contains 00 and such that it is closed under taking unit spheres: all $S(x)$ are in $\mathcal{Q}$ of one dimension less. The Partition monoid actually agrees with this elegant definition:
@@
The class $\mathcal{P}$ of partition graphs has a nice recursive definition. The graphs $\mathcal{P}$ are also characterized number theoretically in that all additive primes in the monoid are zero-dimensional. In the dual Shannon picture, $\hat{\mathcal{P}}$ consists of disjoint union of complete graphs. If one of the additive factors is $1$, it is contractible, meaning that there exists a vertex (actually one can take any vertex), such that $G\setminus\{x\}$ is contractible. The cohomology in general however is not trivial. For the utility graph $K_{(3,3)}$ for example, the Betti vector is $(1,4)$. In general, the Betti vector of $G\in\mathcal{P}$ is $(1,0,\dots,\prod_{j=0}^{k}(n_{j}-1))$ for $k>0$ and $n_{0}$ for $k=0$. In particular, for cross polypes $p=(2,2,\dots,2)$, it is $(1,0,\dots,1)$. The inductive dimension of $G\in\mathcal{P}$ is defined as the average of the dimensions of the unit spheres is equal the maximal dimension. They are small graphs of diameter $1$ or $2$ and have the property that given any finite set of points, then the intersection of all these unit spheres is again in the class. We should think of elements in $\mathcal{P}$ of dimension $k$ therefore as generalized $k$-dimensional points. Complete graphs $K_{k+1}=K_{(1,\dots,1)}$ are just a special case. The intersection of $\mathcal{P}$ and spheres $\mathcal{S}$ is the set of cross polytopes, where every additive prime factor is a zero-dimensional sphere $K_{(2)}=\overline{K_{2}}$. In short, elements in $\mathcal{P}$ can serve as $k$-dimensional points.
@@ 3. Code
RFunction[G_,P_]:=Module[{},R[x_]:=x->RandomChoice[Range[L[Flatten[P]]]];Map[R,Ver[G]]];

### 02 Feb 2024 v2

@@ Abstract.
If $f$ maps a discrete $d$-manifold $G$ onto a $(k+1)$-partite complex $P$ then $H(G,f,P)$, the set of simplices $x$ in $G$ such that $f(x)$ contains at least one facet in $P$ is either a $(d-k)$-manifold or empty.
@@ 1. The theorem
If $f$ maps a $d$-manifold $G$ P-onto a $k$-complex $P\in\mathcal{P}$, then $H(G,f,P)$ is a $(d-k)$-manifold or empty.
@@
Theorem (1) can be seen as a discrete inverse function theorem. It is the discrete analog of the fact that if $f:M\to N$ is a smooth map from a d-manifold $M$ to a $k$-manifold $N$ and $p$ is a regular value in $N$, then $f^{-1}(p)$ is a $d-k$-manifold. The discrete Sard result [8] is a special case, where we used a particular partition complex (topology) on the finite target set $\{-1,1\}^{k}$ of ${\rm sign}(f-c)$ given by a map $f:G\to\mathbb{R}^{k}$. Theorem (1) implies in particular that if $f:G\to y=\{0,\dots,k\}$ is surjective, then the set of simplices $x$ with $f(x)=y$ is a $(d-k)$-manifold or empty. In [6], we looked maps $f:G\to\mathbb{R}$ which leads for $c$ not in $f(G)$ to $x\to{\rm sign}(x)\in\{-1,1\}$. If $f:G\to\mathbb{R}$ has no root and takes both positive and negative values, then $\{f=0\}$ is always a hyper-surface. This was already magic in the very special case $P=K_{2}$.
Figure 1. For Sard $G\to\mathbb{R}$, where studying $\{f=c\}$ is determined by the map ${\rm sign}(f-c)$ onto $\{(-1,1)\}$, the complex $K_{2}=K_{(1,1)}$ is the only relevant choice [6]. For Sard $G\to\mathbb{R}^{2}$, where ${\rm sign}(f-c)$ map into $\{(-1,1),(1,1),(-1,-1),(-1,-1)\}$, we needed to pick the unique 2-dim partition complex $P=K_{(1,1,2)}$ that exists for $n=4$ [8]. For $G\to\mathbb{R}^{3}$, the map ${\rm sign}(f-c)$ targets $\{-1,1\}^{3}$ which has 8 points. The 3-dimensional partitions of $n=8$ are $(5,1,1,1),(4,2,1,1),(3,3,1,1),(3,2,2,1),(2,2,2,2)$. In [8], we picked $p=(1,1,1,5)$ by forced the target signs to be $(-1,1,1),(1,-1,1),(1,1,-1)$ and requiring to reach a 4th point. This complex $P$ is displayed to the right. For maps into $\mathbb{R}^{m}$ we chose $p=(1,1,\dots,1,2^{m}-m)$ in [8].
@@
An interesting case is when the target complex is $P=K_{(2,2,2,2)}$ which is a 3-sphere, the 3-dimensional cross polytop, one of the regular polytops in ambient dimension $4$. We can implement the quaternion group $Q$ on $P$, as we can identify the $8$ points in $P$ with $1,-1,i,-i,j,-j,k,-k$. In other words, a map $f$ from $G$ to the quaternion group $Q$ which not have an Abelian image, produces a co-dimension $3$ manifold $H(G,f,P)$, where $P$ is the 3-sphere simplicial complex structure on $Q$ coming from the partition $p=(2,2,2,2)$. In particular, any map from a 4-manifold to $Q$ with this property produces a co-dimension $3$ manifold $H(G,f,P)$ which must be a finite collection of 1-dimensional cycle graphs.
@@
If $G$ is a $d$-manifold and $f:G\to\{0,1,\dots,k\}$ is surjective then $H$ is a $(d-k)$-manifold or empty.
@@
The empty case is rare. For the smallest 3-sphere $G=K_{(2,2,2,2)}$ and the triangle $P=K_{(1,1,1)}$, there are $5796$ surjective maps. There are only 24 cases, where $H$ is empty. They are the maps to $\{1,2,3\}$ where only one number appears more than once.
1.7.
If $P$ is a $k$-dimensional partition complex and $y$ is a facet in $P$ which is in the image of $f$, then $f^{-1}(y)$ is a $(d-k)$-manifold with boundary or empty. This allows us to construct without much effort manifolds with boundary or pictures of cobordism examples: if the boundary of the manifold is split into two sets $A,B$, then $A,B$ are by definition cobordant. In illustrations done before, like for the expository article [5] in 2012, we needed to construct by hand the ”hose manifold”. We later would use triangulations which numerical methods would provide when parametrizing surfaces $\vec{r}(u,v)$ where $(u,v)\subset\mathbb{R}^{2}$ is a closed region in the parameter plane. The code provided in the code section is shorter than what one needs to get in order to rely on numerical schemes used by computer algebra systems for drawing surfaces. The code works for manifolds of arbitrary dimension.
@@
Every $(d-k)$-manifold $H(G,f,P)$ is the union of patches $f^{-1}(y)$, each of which is a $(d-k)$-manifold with boundary.
@@ 2. Partitions
Figure 4. All 15 integer partitions for $n=7$ and the corresponding 15 complexes. We used Cuisenaire tools in first grade [2] were partitions were studied for the competition [4]. As far as I know, this was the first use of such a pedagogical tool for research rather than for teaching. Beside known results like the Euler pentagonal theorem, unusual results appeared like like $p(n)=\sum_{k=1}^{n}\sigma(k)p(n-k)/n$, where $\sigma(k)$ is the total number of prime factors of $k$. This can be seen by looking at $np(n)$ as the area of the Cuisenaire rectangle listing all the partitions.
@@
Figure 6. A 2-manifold obtained by taking the target complex $P=K_{(2,2,3)}$ and $G=I\oplus C_{13}$, where $I$ is the icosahedron complex. While the graph $G$ has only $12+13=25$ vertices, the graph $H(G,f,P)$ is much larger than $G$ as it is a sub-graph of the Barycentric refinement of $G$.
@@
We now look at a geometric characterization of the partition monoid $(\mathcal{P},\oplus)$. Let $\mathcal{Q}$ be the set of graphs which have the property that it contains 00 and such that it is closed under taking unit spheres: all $S(x)$ are in $\mathcal{Q}$ of one dimension less. The Partition monoid actually agrees with this elegant definition:
@@
The class $\mathcal{P}$ of partition graphs has a nice recursive definition. The graphs $\mathcal{P}$ are also characterized number theoretically in that all additive primes in the monoid are zero-dimensional. In the dual Shannon picture, $\hat{\mathcal{P}}$ consists of disjoint union of complete graphs. If one of the additive factors is $1$, it is contractible, meaning that there exists a vertex (actually one can take any vertex), such that $G\setminus\{x\}$ is contractible. The cohomology in general however is not trivial. For the utility graph $K_{(3,3)}$ for example, the Betti vector is $(1,4)$. In general, the Betti vector of $G\in\mathcal{P}$ is $(1,0,\dots,\prod_{j=0}^{k}(n_{j}-1))$ for $k>0$ and $n_{0}$ for $k=0$. In particular, for cross polytopes $p=(2,2,\dots,2)$, it is $(1,0,\dots,1)$. The inductive dimension of $G\in\mathcal{P}$ is defined as the average of the dimensions of the unit spheres is equal the maximal dimension. They are small graphs of diameter $1$ or $2$ and have the property that given any finite set of points, then the intersection of all these unit spheres is again in the class. We should think of elements in $\mathcal{P}$ of dimension $k$ therefore as generalized $k$-dimensional points. Complete graphs $K_{k+1}=K_{(1,\dots,1)}$ are just a special case. The intersection of $\mathcal{P}$ and spheres $\mathcal{S}$ is the set of cross polytopes, where every additive prime factor is a zero-dimensional sphere $K_{(2)}=\overline{K_{2}}$. In short, elements in $\mathcal{P}$ can serve as $k$-dimensional points.
@@ 3. Code
RFunction[G_,P_]:=Module[{},R[x_]:=x->RandomChoice[Range[L[Ver[P]]]];Map[R,Ver[G]]];
@@
4. Statistics
4.1.
The following code generates all surjective maps and analyzes in each case the topology. This works only for very small host manifolds. To see the topological nature of the manifolds, we look at the cohomology in the form of Betti vectors $(b_{0},b_{1},\dots b_{d})$ which gives the dimensions $b_{k}$ of the harmonic $k$-forms of the complex. This in particular gives the number of connectivity components $b_{0}$ or the genus $b_{1}/2$, which is for surfaces the “number of holes”.
sig[x_]:=Signature[x]; nu[A_]:=If[A=={},0,L[NullSpace[A]]];
F[G_]:=Module[{l=Map[L,G]},If[G=={},{},Table[Sum[If[l[[j]]==k,1,0],{j,L[l]}],{k,Max[l]}]]];
sig[x_,y_]:=If[SubsetQ[x,y]&&(L[x]==L[y]+1),sig[Prepend[y,Complement[x,y][[1]]]]*sig[x],0];
Dirac[G_]:=Module[{f=F[G],b,d,n=L[G]},b=Prepend[Table[Sum[f[[l]],{l,k}],{k,L[f]}],0];
 d=Table[sig[G[[i]],G[[j]]],{i,n},{j,n}]; {d+Transpose[d],b}];
Beltrami[G_]:= Module[{B=Dirac[G][[1]]},B.B];
Hodge[G_]:=Module[{Q,b,H},{Q,b}=Dirac[G];H=Q.Q;Table[Table[H[[b[[k]]+i,b[[k]]+j]],
 {i,b[[k+1]]-b[[k]]},{j,b[[k+1]]-b[[k]]}],{k,L[b]-1}]];
Betti[s_]:=Module[{G},If[GraphQ[s],G=Whitney[s],G=s];Map[nu,Hodge[G]]];
ComputeAllBetti[G_,P_]:=Module[{V=Ver[P],Fa=Facets[P],n=Length[Ver[G]]},
  A=Tuples[V,n]; Surjectiv=Select[A,(L[Union[#]]==L[V]) &];  Print[Length[Surjectiv]];
  Phi[g_]:=Module[{f=Table[k->g[[k]],{k,n}]},H=AbstractSurface[G,f,Fa];Length[H]];
  AllSurfaces=Map[Phi,Surjectiv]; B=Flatten[Position[AllSurfaces,0]];
  EmptyCases=Table[Surjectiv[[B[[k]]]],{k,Length[B]}];
  NonEmptyCases=Complement[Surjectiv,EmptyCases];
  Psi[g_]:=Module[{f=Table[k->g[[k]],{k,n}]},H=AbstractSurface[G,f,Fa];Betti[H]];
  AllBetti=Map[Psi,NonEmptyCases]];
S3 = ZeroJoin[{2,2,2,2}]; S4 = ZeroJoin[{2,2,2,2,2}]; S5 = ZeroJoin[{2,2,2,2,2,2}];
RP3=Generate[{{1,2,3,4},{1,2,3,5},{1,2,4,6},{1,2,5,6},{1,3,4,7},{1,3,5,7},{1,4,6,7},
{1,5,6,7},{2,3,4,8},{2,3,5,9},{2,3,8,9},{2,4,6,10},{2,4,8,10},{2,5,6,11},{2,5,9,11},
{2,6,10,11},{2,7,8,9},{2,7,8,10},{2,7,9,11},{2,7,10,11},{3,4,7,11},{3,4,8,11},
{3,5,7,10},{3,5,9,10},{3,6,8,9}, {3,6,8,11},{3,6,9,10},{3,6,10,11},{3,7,10,11},
{4,5,8,10},{4,5,8,11},{4,5,9,10},{4,5,9,11},{4,6,7,9},{4,6,9,10},{4,7,9,11},{5,6,7,8},
{5,6,8,11},{5,7,8,10},{6,7,8,9}}];
CP2=Generate[{{1,2,3,4,5},{1,2,3,4,7},{1,2,3,5,8},{1,2,3,7,8},{1,2,4,5,6},
  {1,2,4,6,7},{1,2,5,6,8},{1,2,6,7,9},{1,2,6,8,9},{1,2,7,8,9},{1,3,4,5,9},{1,3,4,7,8},
  {1,3,4,8,9},{1,3,5,6,8},{1,3,5,6,9},{1,3,6,8,9},{1,4,5,6,7},{1,4,5,7,9},{1,4,7,8,9},
  {1,5,6,7,9},{2,3,4,5,9},{2,3,4,6,7},{2,3,4,6,9},{2,3,5,7,8},{2,3,5,7,9},{2,3,6,7,9},
  {2,4,5,6,8},{2,4,5,8,9},{2,4,6,8,9},{2,5,7,8,9},{3,4,6,7,8},{3,4,6,8,9},{3,5,6,7,8},
  {3,5,6,7,9},{4,5,6,7,8},{4,5,7,8,9}}];
S2xS2=Generate[{{1,2,3,4,6},{1,2,3,4,7},{1,2,3,6,9},{1,2,3,7,9},{1,2,4,5,8},{1,2,4,5,9},
  {1,2,4,6,8},{1,2,4,7,9},{1,2,5,6,8},{1,2,5,6,9},{1,3,4,6,7},{1,3,5,6,7},{1,3,5,6,9},
  {1,3,5,7,10},{1,3,5,9,11},{1,3,5,10,11},{1,3,7,9,10},{1,3,9,10,11},{1,4,5,8,10},
  {1,4,5,9,11},{1,4,5,10,11},{1,4,6,7,11},{1,4,6,8,10},{1,4,6,10,11},{1,4,7,9,11},
  {1,5,6,7,8},{1,5,7,8,10},{1,6,7,8,11},{1,6,8,10,11},{1,7,8,10,11},{1,7,9,10,11},
  {2,3,4,6,8},{2,3,4,7,8},{2,3,5,7,10},{2,3,5,7,11},{2,3,5,10,11},{2,3,6,8,10},{2,3,6,9,10},
  {2,3,7,8,11},{2,3,7,9,10},{2,3,8,10,11},{2,4,5,8,9},{2,4,7,8,9},{2,5,6,8,11},{2,5,6,9,10},
  {2,5,6,10,11},{2,5,7,8,9},{2,5,7,8,11},{2,5,7,9,10},{2,6,8,10,11},{3,4,6,7,11},{3,4,6,8,10},
  {3,4,6,9,10},{3,4,6,9,11},{3,4,7,8,11},{3,4,8,9,10},{3,4,8,9,11},{3,5,6,7,11},{3,5,6,9,11},
  {3,8,9,10,11},{4,5,6,9,10},{4,5,6,9,11},{4,5,6,10,11},{4,5,8,9,10},{4,7,8,9,11},{5,6,7,8,11},
  {5,7,8,9,10},{7,8,9,10,11}}];
P=ZeroJoin[{1,1,1}];A1=ComputeAllBetti[S3,P];   Save[”allbetti_2222_111.txt”, A1];
P=ZeroJoin[{1,1,1}];A2=ComputeAllBetti[S4,P];   Save[”allbetti_22222_111.txt”,A2];
P=ZeroJoin[{1,1}];  A3=ComputeAllBetti[S4,P];   Save[”allbetti_22222_11.txt”, A3];
P=ZeroJoin[{1,1}];  A4=ComputeAllBetti[RP3,P];  Save[”allbetti_rp3_11.txt”,   A4];
P=ZeroJoin[{1,1,1}];A5=ComputeAllBetti[CP2,P];  Save[”allbetti_cp2_111.txt”,  A5];
P=ZeroJoin[{1,1}];  A6=ComputeAllBetti[CP2,P];  Save[”allbetti_cp2_11.txt”,   A6];
P=ZeroJoin[{1,1}];  A7=ComputeAllBetti[S2xS2,P];Save[”allbetti_s2xs2_11.txt”, A7];
P=ZeroJoin[{1,1,1}];A8=ComputeAllBetti[S2xS2,P];Save[”allbetti_s2xs2_111.txt”,A8];
P=ZeroJoin[{1,2}];  A9=ComputeAllBetti[S3,P];   Save[”allbetti_2222_12.txt”,  A9];
P=ZeroJoin[{1,3}];  A10=ComputeAllBetti[S3,P];  Save[”allbetti_2222_13.txt”, A10];
4.2.
Example 1: In the list of all co-dimension 2 curves in the smallest 3 sphere $K_{(2,2,2,2)}$. There were 4848 single circles,888 cases with 2 circles and 36 cases with 4 circles. There were 5772 surjective maps from $G$ to $K_{3}$ which produced a non-empty surfaces. For only 24 cases, the resulting manifold is empty.
4.3.
Example 2: Take the smallest 4 sphere $K_{(2,2,2,2,2)}$. There were 55980 surjective maps from $G$ to $K_{3}$. This produces co-dimension 2 surfaces. Only 30 surjective function sproduced an empty surface. The other 55950 produced either 2-spheres (50160), 2 disjoint 2-spheres (3960) or 4 disjoint 2-spheres (60 cases) or single $2$-tori (1680 cases) or 2 disjoint $2$-tori (90 cases).
4.4.
Example 3: For all co-dimension $1$ surfaces in the 4-sphere $G=K_{(2,2,2,2,2)}$, we see $992$ 3-spheres, $20$ cylinders $S^{2}\times S^{1}$ and $10$ pairs of disjoint sphere. There were $1022$ surjective maps from $G$ to $K_{2}$, so that 30 produced empty sets.
4.5.
Example 4: All co-dimension 1 surfaces in a small projective 3-manifold $\mathbb{RP}^{3}$ produce $22046=2^{12}-$ 2 possible surjective maps to $K_{2}$. All of them give surfaces. There are $548$ spheres, $1394$ tori, $68$ genus 2 surfaces, $8$ genus 3 surfaces, $20$ pairs of spheres, $8$ pairs with one sphere and one torus and 2 pairs, where one is a $2$-torus and one a genus 2 surface.
4.6.
Example 5: When looking at co-dimension 2 surfaces in a small implementation of the 4-manifold $\mathbb{CP}^{2}$, we count $8748$ 2-spheres and $9402$ tori.
4.7.
Example 6: Among all co-dimension 1 surfaces in a small implementation of $\mathbb{CP}^{2}$ we count $511$ 3-spheres.
4.8.
Example 7 For all hyper-surfaces defined by $K_{2}$ in the small 4-manifold $\mathbb{S}^{2}\times\mathbb{S}^{2}$ defined by surjective to $K_{2}$ the nuumber of $3$-manifolds is 2046. There were 1446 3-spheres with Betti vector $(1,0,0,1)$ and 600 generalized tori $S^{2}\times S^{1}$ with Betti vector $(1,1,1,1)$.
4.9.
Example 8 looks for co-dimension 2 surfaces in a small $G=S^{2}\times S^{2}$ using the triangle $P=K_{3}$. There are $42288$ two-spheres, $71022$ genus 1 surfaces ($2$-tori), $48210$ genus 2 surfaces, $7788$ genus 3 surfaces, $660$ genus 4 surfaces, $78$ genus 5 surfaces and 960 disconnected pairs of 2-spheres. In total, the cohomology of $171006$ surfaces was computed.
4.10.
Example 9 For all hyper-surfaces in the small 3 sphere $K_{(2,2,2,2)}$ with $P=K_{(1,2)}$ we have $5496$ surfaces, where $5456$ were 2-spheres, $84$ were $2$-tori and $256$ were double spheres.
4.11.
Example 10 For all hyper-surfaces in the small 3 sphere $G=K_{(2,2,2,2)}$ with $P=K_{(1,3)}$ we have $40824$ cases in total, no empty surfaces, $38448$ spheres, $216$ tori and $2160$ double spheres. In this case we look at all maps from $G$ to $\{1,2,3,4\}$ because $P$ has $4$ vertices. This requires to look at much more maps. An interesting question is how large the total number of hyper-surfaces $\bigcup_{P}\bigcup_{f}H(G,f,P)$ can become. As they are all sub-graphs of the Barycentric refinement $G_{1}$ of $G$, there are only finitely many. Many $H(G,f,P)$ are are actually the same graph.
4.12.
As the gallery in the next section illustrates, much more interesting sub-manifolds are possible for larger host manifolds $G$. But the number of possible maps is too large for an exhaustive list. We would have to use Monte-Carlo simulations to investigate what manifolds typically occur.